# TIPO: Text to Image with Text Presampling for Prompt Optimization

### v1

@@ Appendix F Topic Distribution Visualization
To analyze the meaningfulness and diversity of topics within text inference results, we implemented a multi-stage topic modeling and clustering methodology. Initially, we employed Latent Dirichlet Allocation (LDA) to extract varying numbers of topics—specifically 20, 30, 50, and 100—from the corpus [8]. This approach aimed to identify representative words that are significant within the dataset while filtering out common stop words, ensuring that nonsensical or irrelevant terms are excluded from topic classification.
By employing multi-level topic analysis, we empirically assessed whether the resulting topics were sufficiently large and diverse. This iterative process mitigates potential challenges such as substantial topic overlap, which can diminish distinctiveness when extracting a large number of topics (Stevens et al., 2012).
To address the potential overlap and further assess the diversity and meaningfulness of the topics, we performed a secondary clustering. We grouped the initially extracted topics into five major clusters using k-means clustering and evaluated the clustering performance by calculating the inertia (Sum of Squared Distances) [23], shown in Table 4, 5, and 6. Since the topics have already been filtered for meaningful content, a higher inertia value indicates greater diversity among the clusters, reflecting a broader range of valid and meaningful topics across the dataset. This two-tiered approach allows for a more nuanced analysis of topic diversity and ensures the robustness of the topic modeling against meaningless word groupings.
Table 4: Inertia for COYO-Dataset inference, higher is better
@@
Table 5: Inertia for GBC-Dataset inference, higher is better
@@
Table 6: Inertia for Scenery extend inference, higher is better

### v2

@@ Appendix F Topic Distribution Visualization
Latent Dirichlet Allocation (LDA) [9] is a generative probabilistic model for topic modeling [31], which assumes that each document is a mixture of topics, with each topic represented by a distribution over words. LDA uncovers hidden thematic structures by analyzing word co-occurrence patterns, while methods like TF-IDF and TextRank [42] enhance its ability to extract meaningful insights from large textual datasets. We implemented a multi-stage topic modeling and clustering methodology using LDA to extract varying numbers of topics (20, 30, 50, and 100) from the corpus. This approach focuses on identifying significant representative words while filtering out stop words and irrelevant terms to ensure meaningful topic classification.
By employing multi-level topic analysis, we empirically assessed whether the resulting topics were sufficiently large and diverse. This iterative process mitigates potential challenges such as substantial topic overlap, which can diminish distinctiveness when extracting a large number of topics [59].
To address the potential overlap and further assess the diversity and meaningfulness of the topics, we performed a secondary clustering [66]. We grouped the initially extracted topics into five major clusters using k-means clustering and evaluated the clustering performance by calculating the inertia (Sum of Squared Distances) [27], shown in Table 7, 8, and 9. Since the topics have already been filtered for meaningful content, a higher inertia value indicates greater diversity among the clusters, reflecting a broader range of valid and meaningful topics across the dataset. This two-tiered approach allows for a more nuanced analysis of topic diversity and ensures the robustness of the topic modeling against meaningless word groupings.
@@
Table 7: Inertia for COYO-Dataset inference, higher is better
@@
Table 8: Inertia for GBC-Dataset inference, higher is better
@@
Table 9: Inertia for Scenery extend inference, higher is better
@@
Figure 19: This visualization represents a filtered subset of posts from the Danbooru2023 dataset, centered on the ’scenery’ tag. The network graph is an ego network (depth 1), meaning it includes only nodes directly connected to the ’scenery’ tag. To refine the data and focus on meaningful associations, uncommon tags with fewer than 10 occurrences were excluded. The analysis, conducted using Gephi, focuses on nodes with a degree greater than 600 to highlight critical components. Nodes are color-coded by modularity class by Fast Unfolding Algorithm [10], revealing clusters of closely associated tags. Node size reflects Eigenvector Centrality[11], emphasizing tags that are both highly connected and influential within their network.